OpenStudy (anonymous):

Choose the correct simplification of the expression (4a4b)2. 16a8b 16a8b2 16a6b2 8a8b2

OpenStudy (phi):

you have to put in ^ because it is not clear what is an exponent is this \[(4a^4b)^2 \] ? if so write it (4 a^4 b)^2

OpenStudy (phi):

\[ (4 a^4 b)^2 = 4 a^4 b 4 a^4 b \] in other words, multiplied by itself twice

OpenStudy (anonymous):

whats the answer

OpenStudy (phi):

First, notice you have 4*4 so the number part of the answer is 16

OpenStudy (phi):

what is a^4 * a^4 ? (remember a^4 is short-hand for a*a*a*a, so a^4 * a^4 = a*a*a*a* a*a*a*a do you know how to write a*a*a*a* a*a*a*a using the short-hand ?

OpenStudy (anonymous):

is it A B C or D